6. Two cars are traveling along a straight road. Car (A) maintains a constant speed of 90.0 km/h; car (B) maintains a constant speed of 104 km/h. At t car (B) is 43.0 km bchind car A. How far will car (A) travel from t- 0 bcforc it is overtaken by car (B)? 0, x 289 km
Answer #1

Car B is gaining on car A at 104-90 = 14 km/hr

The time for B to overtake A is 43/14 = 3.07 hr

In that time car A will travel 3.07 * 90 = 276.3 km